OpenStudy (anonymous):

Find the vertex and line of symmetry y=-x^2+10x-34

sam (.sam.):

Vertex=complete the square (remember)

OpenStudy (anonymous):

vertex = -b/2a

sam (.sam.):

-x^2+10x-34 -(x^2-10x)-34 -(x-5)^2-25(-1)-34 -(x-5)^2-9

sam (.sam.):

Vertex= (5,-9) Symmetry at x=5

sam (.sam.):

@jazy well complete the square get both the solutions at a time

OpenStudy (saifoo.khan):

and it will be a maximum point.

OpenStudy (anonymous):

where did you get the -25(-1)

OpenStudy (anonymous):

That's true I guess, but I'm only in Algebra 1 so I only know the basic ways of solving these.

OpenStudy (anonymous):

ya they are a little confusing

sam (.sam.):

@babygirl81 that's completing the square with fast method, I'll teach you, For example: x^2+bx+c You divide 2 for 'b' Then you get (x+b/2)^2 Then square the b/2 and put it outside but its always negative (x+b/2)^2-(b/2)^2+c

OpenStudy (anonymous):

oh ok so then you multiply by -1 to make it positive, is that why you did that?

OpenStudy (anonymous):

This is how I would have solved it: x = -b/2a x = -10/2(-1) btw I think this is what @.sam. put because -x = -1x x = -10/-2 x = 5 y=-1(5)^2+10(5)-34 y= -1(25) + 50 - 34 y = -25 + 50 - 34 y = -9 (5,-9) Line of symmetry at x=5 A LOT LONGER lol....but I'm just learning now:) The other method seems pretty short and simple though:)

OpenStudy (anonymous):

oh ok between both of those I get it now that where the -1 came from that is a.

OpenStudy (anonymous):

jazy did the substitution method and sam did the quick method the substitution method is a good way to learn how to do it thats how i did it before but i forgot that was a couple of chapters ago, just studying for the final